OpenStudy (anonymous):

An equation in point slope form, standard form, and slope-intercept form. for the line that passes through (0, -2), (3, 2). Will fan and medal

563blackghost (563blackghost):

First we would find the slope so use the slope formula... \[\frac{ y2-y1 }{ x2-x1 }\]

563blackghost (563blackghost):

We input the points and subtract and would get the slope...

OpenStudy (anonymous):

Yes. Ok so the slope would be 1 and 1/3rd.

563blackghost (563blackghost):

Incorrect.....\[\frac{ 2--2 }{ 3-0 }\] Remember that when subtracting a negative that you are practically adding so it would turn to 2 + 2....\[\frac{ 2+2 }{ 3-0 }\]

OpenStudy (anonymous):

3-0 = 3... 4/3 = 1.33333 so 1 and 1/3 or 4/3 if you want to call it that.

563blackghost (563blackghost):

Yup it would be 4/3 ^^ Now lets input that in slope-intercept form...\[y=mx+b\] So m is the slope and b is the y-intercept...now the y-intercept is where we know y while x is 0 so (0,#)....do we know the y-intercept?

OpenStudy (anonymous):

Wouldn't the Y intercept just be either, -2 or 2?

563blackghost (563blackghost):

it would be -2 xD so how would the equation look like?

OpenStudy (anonymous):

-2 = 4/3 * 0 + B

563blackghost (563blackghost):

Incorrect.... remember x will stay x and the y-intercept goes where b is...

OpenStudy (anonymous):

Uh, ok so B is -2?

563blackghost (563blackghost):

yup ^^

OpenStudy (anonymous):

So, y = 4/3 * 0 + -2

563blackghost (563blackghost):

Correct! so...\[y=\frac{ 4 }{ 3 }x-2\]

563blackghost (563blackghost):

Now lets get point slope form.... Now for point slope we must know the first coordinates and the slope which we know of so we would input it into the equation....\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

ok, what we just did was slope intercept form right

563blackghost (563blackghost):

Correct ^^

OpenStudy (anonymous):

Ok.. So, can you explain the point slope a bit more thoroughly?

563blackghost (563blackghost):